# ReliableMath: Benchmark of Reliable Mathematical Reasoning on Large Language Models

### v1

@@ Abstract
Although demonstrating remarkable performance on reasoning tasks, Large Language Models (LLMs) still tend to fabricate unreliable responses when confronted with problems that are unsolvable or beyond their capability, severely undermining the reliability. Prior studies of LLM reliability have primarily focused on knowledge tasks to identify unanswerable questions, while mathematical reasoning tasks have remained unexplored due to the dearth of unsolvable math problems. To systematically investigate LLM reliability in mathematical reasoning tasks, we formulate the reliability evaluation for both solvable and unsolvable problems. We then develop a ReliableMath dataset which incorporates open-source solvable problems and high-quality unsolvable problems synthesized by our proposed construction workflow with human evaluations. Experiments are conducted on various LLMs with several key findings uncovered. LLMs fail to directly identify unsolvable problems and always generate fabricated responses. When instructing LLMs to indicate unsolvability using a reliable prompt, the reliability of larger-sized LLMs remains on solvable problems, but notably improves on unsolvable problems yet still falls short of solvable problems. However, small LLMs rarely show any progress despite employing reliable prompts. Therefore, we further propose an alignment strategy to enhance small LLMs’ reliability, which can significantly improve LLM reliability performances on both in-domain and out-of-domain tasks.
@@ 1 Introduction
Large Language Models (LLMs), such as DeepSeek-R1 deepseekai2025deepseekr1incentivizingreasoningcapability and OpenAI-o1 (openai2024o1, ), have exhibited impressive capabilities in reasoning tasks li202512surveyreasoning ; chen2025reasoningerasurveylong . However, when confronted with problems that are intrinsically unsolvable or beyond their capability scopes (yin-etal-2023-large, ; amayuelas-etal-2024-knowledge, ; yang2024alignment, ; liu2023trustworthy, ; li2024surveyhonestylargelanguage, ; xue2024ualignleveraginguncertaintyestimations, ), LLMs may still attempt to fabricate reasoning steps to provide plausible but misleading answers to users, as illustrated in Fig. 1 (a). This behavior results in substantial factually incorrect or nonsensical content, also referred to as the “hallucination” issue (Huang_2025, ; 10.1145/3571730, ), thereby potentially undermining LLMs’ reliability.
The reliability of LLMs necessitates generating factually correct, informative, and trustworthy content (liu2023trustworthy, ). Prior research on LLM reliability has predominantly focused on knowledge-based tasks, which identify LLMs’ known and unknown/answerable questions (yin-etal-2023-large, ; amayuelas-etal-2024-knowledge, ; zheng2025enhancingllmreliabilityexplicit, ; li2024surveyhonestylargelanguage, ), lacking attention on typical reasoning tasks like mathematical reasoning (tang2024mathscalescalinginstructiontuning, ; liu2025acemathadvancingfrontiermath, ). Specifically, mathematical reasoning problems can be divided into solvable and unsolvable problems depending on whether a reasonable solution that satisfies all requirements of the problem exists. Determining the solvability of problems or whether LLMs can solve requires thoughtful reasoning step by step, intensifying the challenges of reliability evaluation on reasoning tasks. Additionally, existing mathematical benchmarks exclusively concentrate on solvable problems (tang2024mathscalescalinginstructiontuning, ; liu2025acemathadvancingfrontiermath, ), while unsolvable/unknown datasets are confined to knowledge tasks (yin-etal-2023-large, ; amayuelas-etal-2024-knowledge, ), leaving a scarcity of unsolvable math problems for study.
Accordingly, this work systematically investigates LLM reliability on mathematical reasoning tasks. We formulate LLM reliability evaluation on both solvable and unsolvable mathematical problems. As in Fig. 1 (b), a reliable LLM on mathematical reasoning as one that should correctly answer solvable problems or explicitly identify the unsolvability of unsolvable problems after step-by-step reasoning, which is the most preferable and considered as successful as in Fig. 1 (c). If the problem falls beyond the LLMs’ capacity to determine the solvability, a suboptimal response would be refusing to answer (tang2024mathscalescalinginstructiontuning, ; xu2024rejection, ; yang2024alignment, ; xue2024ualignleveraginguncertaintyestimations, ) to prevent misleading users. All other responses are considered as failed.
@@
We then develop a ReliableMath dataset, including both solvable and unsolvable data. Solvable math problems are collected from open-source datasets like AIME (aimo2024aime, ). Unsolvable problems are synthesized using our proposed three-stage data construction workflow. First, we rewrite the solvable problems to render them unsolvable. Then, we verify the rewritten problems using LLMs to judge whether they are unsolvable, and filter out unqualified data. Finally, we employ human experts to check and maintain high-quality unsolvable problems to constitute the ReliableMath dataset.
@@
Experiments are conducted on ReliableMath dataset on a series of reasoning LLMs (li202512surveyreasoning, ; chen2025reasoningerasurveylong, ) like DeepSeek-R1 (deepseekai2025deepseekr1incentivizingreasoningcapability, ) and instruction LLMs like GPT-4o (openai2024gpt4o, ). Our analysis yields several key findings. LLMs cannot directly identify the unsolvability of problems and will attempt to fabricate reasoning steps with considerable tokens, severely diminishing the reliability. By employing a reliable prompt that can instruct LLMs to indicate the unsolvability, the reliability of large-sized LLMs remains on solvable problems, but significantly improves on unsolvable cases, although the improved reliability on unsolvable problems still falls short of solvable data. After using reliable prompts, larger-size reasoning LLMs generally outperform instruction LLMs in reliability. However, both small reasoning and instruction LLMs barely exhibit any gains on unsolvable problems.
@@
Furthermore, we employ an alignment strategy to boost small LLMs’ reliability. Specifically, we adopt the data construction workflow to obtain unsolvable mathematical training data, and sample reliable responses by rejection sampling (yuan2023scalingrelationshiplearningmathematical, ) for both solvable and unsolvable problems, which are utilized to train small LLMs to align with reliability. Results show that the alignment method can significantly enhance LLM reliability on both in-domain and out-of-domain tasks.
The main contributions of this work are presented as follows:
@@
1) This work is the first to comprehensively investigate LLM reliability on reasoning tasks. We formulate the reliability evaluation on both solvable and unsolvable mathematical problems for LLMs, providing a view into assessing LLM reliability on mathematical reasoning.
2) We develop a dataset including unsolvable math problems obtained by our proposed data construction workflow with human evaluations, addressing the gap of unsolvable mathematical datasets.
@@
3) We conduct experiments on various reasoning and instruction LLMs for reliability evaluation, showcasing several valuable findings to inspire developing more reliable LLMs in future work.
@@
4) We propose an alignment strategy that can effectively enhance LLMs’ reliability on both in-domain and out-of-domain tasks, providing insights for further improvements of LLM reliability.
2 Reliability Evaluation
2.1 Definition and Formulation of Reliability
@@
Reliable LLMs are supposed to generate factual, informative, and trustworthy content (liu2023trustworthy, ). Although specific definitions of reliability are many-sided, a unified perspective is that LLMs should identify what they know and can answer, and what they are unable to answer (li2024surveyhonestylargelanguage, ; NEURIPS2024_0d99a8c0, ; liu2023trustworthy, ). Hence, on knowledge tasks, given a question $\boldsymbol{x}$, ground truth $\boldsymbol{y}$, and an LLM $\mathcal{M}$, identifying known/unknown facts in LLM-generated response $\boldsymbol{\hat{y}}=\mathcal{M}(\boldsymbol{x})$ is widely employed for reliability evaluation (yin-etal-2023-large, ; amayuelas-etal-2024-knowledge, ).
@@
Definition 1 (LLM Reliability on Reasoning)
A reliable LLM $\mathcal{M}$ should be capable of identifying the solvability of problem $\boldsymbol{x}$, and for a solvable question, $\mathcal{M}$ can provide correct reasoning step $\boldsymbol{\hat{r}}$ and answer $\boldsymbol{\hat{y}}$, while for an unsolvable question, $\mathcal{M}$ can explicitly analyze and indicate the unsolvability in $\boldsymbol{\hat{r}}$ and $\boldsymbol{\hat{y}}$. If failing to determine the solvability, a suboptimal choice for $\mathcal{M}$ is to refuse in $\boldsymbol{\hat{y}}$ for both solvable and unsolvable cases.
Successful $\mathcal{S}$ Refused $\mathcal{R}$ Failed $\mathcal{F}$
@@
As depicted in Fig. 1 (b), Definition 1 and Table 1, we formulate the LLM reliability evaluation on reasoning tasks as follows. The questions are categorized along two dimensions — Solvable ($\mathcal{A}$) and Unsolvable ($\mathcal{U}$) — and LLM responses along three dimensions - Successful ($\mathcal{S}$), Refused ($\mathcal{R}$), and Failed ($\mathcal{F}$). A successful $\boldsymbol{\hat{y}}$ exactly matches the ground truth $\boldsymbol{y}$ providing the correct answer for $\boldsymbol{x}\in\mathcal{A}$ or stating the problem is unsolvable for $\boldsymbol{x}\in\mathcal{U}$. Refused responses express “I don’t know” in $\boldsymbol{\hat{y}}$ for both $\mathcal{A}$ and $\mathcal{U}$. All other cases are considered as failed. Since directly evaluating intermediate steps $\boldsymbol{\hat{r}}$ is demanding, our assessment is confined to the final answer $\boldsymbol{\hat{y}}$. For both both $\mathcal{A}$ and $\mathcal{U}$, the preference of $\boldsymbol{\hat{y}}$ is: $\mathcal{S}$$\mathcal{R}$$\mathcal{F}$. This formulation offers an insight into design evaluation metrics and improvement methods, thereby enhancing LLMs’ reliability.
2.2 Reliability Evaluation Metrics
We employ two metrics to evaluate reliability: Precision (Prec.) and Prudence (Prud.). Precision measures the proportion of $\mathcal{S}$ calculated on $\mathcal{A}$ and $\mathcal{U}$ separately, and then averaging as follows.
[…]
Prudence represents the proportion of $\mathcal{R}$, which is also computed on $\mathcal{A}$ and $\mathcal{U}$, and then averaged.
[…]
When assessing LLMs’ reliability, precision is prioritized, and the higher the precision, the better the reliability. When achieving comparable precisions for two LLMs, prudence is then considered, with a higher prudence being preferable, as in Fig. 1 (b), Definition 1 and Table 1.
@@ 3.1 Solvable Data Collection
Mathematical tasks are widely utilized to assess LLMs’ reasoning capabilities (tang2024mathscalescalinginstructiontuning, ; liu2025acemathadvancingfrontiermath, ). Accordingly, this study employs four representative open-source datasets, spanning from high-school-level data MATH (hendrycks2021measuring, ) to competitive college-level problems MinervaMath (minervamath2024, ) and Olympic-level challenges AIME, AMC (aimo2024aime, ; aimo2024amc, ). Full sets of 30 AIME and 83 AMC problems are incorporated, alongside 100 problems randomly sampled from MinervaMath and 100 from MATH respectively, constituting the solvable subset $\mathcal{D}_{a}=\{(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\}_{i=1}^{N}$. More dataset details are provided in Appendix F.1.
@@ 3.2 Unsolvable Data Rewriting Types
Figure 2: Illustrations of (a) one solvable question from AIME aimo2024aime and its two rewritten questions by removing one condition (b) or adding one contradictory condition (c).
Current research on unknown or unsolvable problems has predominantly focused on knowledge-intensive tasks (amayuelas-etal-2024-knowledge, ; yin_large_2023, ), with limited attention to reasoning tasks like mathematics. Therefore, we propose an unsolvable mathematical dataset for this gap. To ensure unsolvable problems under the same distribution of solvable data sources, we construct the unsolvable dataset $\mathcal{D}_{u}$ by rewriting solvable problems in $\mathcal{D}_{a}$ into unsolvable forms. The rewriting process is grounded in the following premise: for a given problem $\boldsymbol{x}_{i}$ containing $k$ necessary mathematical conditions, a unique solution $\boldsymbol{y}_{i}$ is derived by reasoning with all these conditions $\{{\boldsymbol{c}_{i}^{1}},{\boldsymbol{c}_{i}^{2}},\dots,{\boldsymbol{c}_{i}^{k}}\}\in\boldsymbol{x}_{i}$, as illustrated in Fig. 2 (a). Any alteration to one necessary condition may result in the loss of a unique solution (fan2025missingpremiseexacerbatesoverthinking, ). Accordingly, we employ two approaches to synthesize unsolvable problems:
  • Removal: As in Fig. 2 (b), by removing one key mathematical premise ${\boldsymbol{c}_{i}^{2}}$, the original reasoning steps are prevented, thereby rendering the problem unsolvable.
  • Contradiction: As in Fig. 2 (c), introducing a condition ${\boldsymbol{\tilde{c}}_{i}^{1}}$ that contradicts the condition ${\boldsymbol{c}^{1}_{i}}$ creates a logical inconsistency within $\boldsymbol{x}_{i}$, making it impossible to derive a solution.
Simply modifying a condition may sometimes result in false unsolvable behaviors where some rewritten problems have non-unique solutions or yield a solution that differs from $\boldsymbol{y}_{i}$, rather than genuinely unsolvable. Therefore, we further propose a rigorous workflow in Sec. 3.3.

### v2

# ReliableMath: Benchmark of Reliable Mathematical Reasoning for Large Language Models

@@ Abstract
Large Language Models (LLMs) tend to fabricate unreliable responses to problems that are unsolvable or beyond their capabilities, severely undermining the reliability. Previous work has mainly examined reliability on knowledge tasks, leaving math reasoning largely unexplored due to the dearth of unsolvable math problems. To investigate LLM reliability on such math tasks, we formulate the reliability evaluation for both solvable and unsolvable problems. We then develop a ReliableMath dataset including open-source solvable problems and high-quality unsolvable problems synthesized by our designed workflow with expert check. Experiments are conducted on various LLMs with several key findings uncovered: 1) LLMs can occasionally recognize the illogicality of unsolvable problems, but always fail to directly identify the unsolvability instead of fabricating reasonings. 2) When instructing LLMs to critically identify solvability with reliable prompts, the reliability performance of larger-sized LLMs remains on the solvable, and notably improves on unsolvable problems yet still lags behind on solvable cases. 3) Small LLMs rarely show any progress despite employing reliable prompts. Therefore, we further propose an alignment strategy to enable small LLMs to critically think and identify problems, which can significantly improve reliability on both in-domain math and out-of-domain knowledge tasks.
@@ 1 Introduction
Large Language Models (LLMs), such as DeepSeek-R1 deepseekai2025deepseekr1incentivizingreasoningcapability and OpenAI-o1 (openai2024o1), have exhibited impressive capabilities in reasoning tasks li202512surveyreasoning; chen2025reasoningerasurveylong. However, LLMs tend to produce deterministic responses to any questions after pre-training (tian2024finetuning). Hence, as illustrated in Fig. 1 (a), when confronted with problems that are intrinsically unsolvable or beyond their capabilities (yin-etal-2023-large; amayuelas-etal-2024-knowledge; yang2024alignment; liu2023trustworthy; li2024surveyhonestylargelanguage; xue2024ualignleveraginguncertaintyestimations), LLMs may still attempt to fabricate reasoning steps to output plausible but misleading answers, including substantial factually incorrect, illogical, or nonsensical content, also referred to as “hallucination” (Huang_2025; 10.1145/3571730), severely undermining LLM reliability.
@@
LLM reliability necessitates generating factually correct, logical, and informative content (liu2023trustworthy). Prior research assessing LLM reliability has predominantly focused on LLMs’ abilities to identify their known/unknown questions on knowledge-based tasks (yin-etal-2023-large; amayuelas-etal-2024-knowledge; zheng2025enhancingllmreliabilityexplicit; li2024surveyhonestylargelanguage), lacking attention on reasoning tasks like mathematics (tang2024mathscalescalinginstructiontuning; liu2025acemathadvancingfrontiermath). Generally, educational-level mathematical problems may be either solvable or unsolvable, depending on whether a reasonable solution that satisfies all the problem’s requirements exists.  [^1] In a general sense, real-world unsolvable math problems include both illogical, ill-posed, or undecidable problems (e.g. Find a real solution to an equation with no real solutions.) (unsolvable; mathematical) and unsolved open problems (e.g. Hilbert’s problems) (unsolved). Since unsolved problems are limited and public, unsolvable math problems in this work are confined to ill-posed, logically unsolvable problems where no definitive solution exists. Determining the solvability of problems requires thoughtful reasoning step by step, intensifying the challenges of reliability evaluation on reasoning tasks. Additionally, existing mathematical benchmarks exclusively concentrate on solvable problems (tang2024mathscalescalinginstructiontuning; liu2025acemathadvancingfrontiermath), while unanswerable/unknown datasets are confined to knowledge tasks (yin-etal-2023-large; amayuelas-etal-2024-knowledge), leaving a scarcity of high-quality open-source unsolvable mathematical problems, which can be utilized to assess LLM reasoning reliability (ma2025largelanguagemodelsstruggle).
To this end, this work systematically investigates LLM reliability on math reasoning tasks. We formulate the LLM reliability definition on both solvable and unsolvable mathematical problems as follows: a reliable LLM should correctly answer solvable problems or explicitly identify the unsolvability of unsolvable problems after step-by-step reasoning as in Fig. 1 (b), which is the most preferable and considered as successful in Fig. 1 (c). If the problem falls beyond the LLMs’ capacity to answer, a suboptimal response would be refusing to answer (tang2024mathscalescalinginstructiontuning; xu2024rejection; yang2024alignment) to prevent misleading users. All other responses are considered as failures.
@@
We then develop a ReliableMath dataset, including both solvable and unsolvable data. Solvable math problems are collected from open-source datasets at different levels. Unsolvable problems are synthesized using our proposed three-stage data construction workflow including 1) solvable problems are rewritten by removing necessary conditions or incorporating contradictions, 2) rewritten problems are verified whether they are unsolvable using LLMs, and 3) verified problems are checked by human experts and qualified unsolvable problems are maintained to constitute the high-quality ReliableMath dataset. The data synthesis workflow is scalable to generate unsolvable mathematical problems at any difficulty level for reasoning reliability evaluation of LLMs.
@@
Experiments on ReliableMath dataset are conducted on a series of reasoning LLMs (li202512surveyreasoning; chen2025reasoningerasurveylong) like DeepSeek-R1 (deepseekai2025deepseekr1incentivizingreasoningcapability) and instruction LLMs like GPT-4o (openai2024gpt4o). Our analysis yields several key findings: 1) LLMs can sometimes recognize the illogicality of unsolvable problems, but always fail to directly identify the unsolvability of problems instead of fabricating reasoning steps with considerable tokens, severely diminishing the reliability. 2) By employing reliable prompts that instruct LLMs to indicate unsolvability, the reliability of large-sized LLMs persists on solvable problems, but significantly improves on unsolvable cases, although the improved reliability on unsolvable problems still falls short of solvable data. 3) After using reliable prompts, large-sized reasoning LLMs generally outperform instruction LLMs in reliability, while all small LLMs barely exhibit any gains on unsolvable problems.
@@
Furthermore, we employ an alignment strategy to enable small LLMs to critically think and identify the solvability of problems. Specifically, we adopt the data construction workflow to obtain unsolvable mathematical training data, and sample reliable responses by rejection sampling (yuan2023scalingrelationshiplearningmathematical), which are utilized to train small LLMs to align with reliability. Results show that the alignment method can significantly improve reliability on both in-domain (ID) math tasks and out-of-domain (OOD) knowledge tasks, where OOD tasks contain both open-source known questions and real-world unknown questions collected from crowd-source workers (amayuelas-etal-2024-knowledge).
The main contributions of this work are below:
@@
1) We first comprehensively investigate LLM reliability on reasoning tasks and formulate the reliability evaluation on both solvable and unsolvable mathematical problems, providing a view into assessing LLM reasoning reliability.
2) We develop a dataset including unsolvable math problems by our proposed data construction workflow with expert check, which is scalable to synthesize different-level unsolvable math problems for reasoning reliability evaluation of LLMs.
@@
3) We conduct experiments on various LLMs on the constructed ReliableMath dataset, showcasing several valuable findings to inspire developing more reliable LLMs in future work.
@@
4) We propose an alignment strategy to effectively enhance LLMs’ reliability on both in-domain and out-of-domain tasks, providing insights for further improvements of LLM reliability.
2 Definition of Reliability
@@
Reliable LLMs are supposed to generate factual and informative content (liu2023trustworthy). Although specific definitions of reliability are many-sided, a unified perspective is that LLMs should identify what they know and can answer, and what they are unable to answer (li2024surveyhonestylargelanguage; NEURIPS2024_0d99a8c0; liu2023trustworthy). Hence, on knowledge tasks, given a question $\boldsymbol{x}$, ground truth $\boldsymbol{\hat{y}}$, and an LLM $\mathcal{M}$, identifying known/unknown facts in LLM-generated response $\boldsymbol{y}=\mathcal{M}(\boldsymbol{x})$ is widely employed for reliability evaluation (yin-etal-2023-large; amayuelas-etal-2024-knowledge; yang2024alignment).
@@
Definition 1 (LLM Reasoning Reliability)
A reliable LLM $\mathcal{M}$ should be capable of identifying the solvability of problem $\boldsymbol{x}$, and for a solvable question, $\mathcal{M}$ can provide correct reasoning step $\boldsymbol{r}$ and answer $\boldsymbol{y}$, while for an unsolvable question, $\mathcal{M}$ can explicitly analyze and indicate the unsolvability in $\boldsymbol{r}$ and $\boldsymbol{y}$. If failing to determine the solvability, a suboptimal choice for $\mathcal{M}$ is to refuse in $\boldsymbol{y}$ for both solvable and unsolvable cases.
Success $\mathcal{S}$ Refusal $\mathcal{R}$ Failure $\mathcal{F}$
@@
As depicted in Fig. 1 (b), Def. 1 and Table 1, we formulate the LLM reliability on reasoning tasks as follows. The questions are categorized along two dimensions - Solvable ($\mathcal{A}$) and Unsolvable ($\mathcal{U}$) - and LLM responses along three dimensions - Success ($\mathcal{S}$), Refusal ($\mathcal{R}$), and Failure ($\mathcal{F}$). A successful $\boldsymbol{y}$ exactly matches the ground truth $\boldsymbol{\hat{y}}$, which provides the correct answer for $\boldsymbol{x}\in\mathcal{A}$ or stating the problem is unsolvable for $\boldsymbol{x}\in\mathcal{U}$ after step-by-step reasoning. Refused responses express “I don’t know” in $\boldsymbol{y}$ for both $\mathcal{A}$ and $\mathcal{U}$. All other cases are considered as failed. For both $\mathcal{A}$ and $\mathcal{U}$, the preference of $\boldsymbol{y}$ is: $\mathcal{S}$$\mathcal{R}>>\mathcal{F}$. This formulation offers an insight into further designing evaluation metrics and alignment strategies to enhance LLMs’ reliability in the following sections.
@@ 3.1 Solvable Data Collection
Math tasks are widely utilized to assess LLMs’ reasoning capabilities (tang2024mathscalescalinginstructiontuning; liu2025acemathadvancingfrontiermath). Accordingly, this study employs four representative open-source datasets, spanning from high-school-level data MATH (hendrycks2021measuring) to competitive college-level problems MinervaMath (minervamath2024) and Olympic-level challenges AIME, AMC (aimo2024aime; aimo2024amc). Full sets of 30 AIME and 83 AMC problems are incorporated, alongside 100 problems randomly sampled from Minerva and 100 from MATH respectively, constituting the solvable subset $\mathcal{D}_{a}=\{(\boldsymbol{x}_{i},\boldsymbol{\hat{y}}_{i})\}_{i=1}^{N}$. Dataset details are provided in Supplement 9.
@@ 3.2 Unsolvable Data Rewriting Types
Figure 2: Illustrations of (a) a solvable question from AIME aimo2024aime and two rewritten schemes by removing one condition (b) or adding one contradictory condition (c).
Current research on unanswerable problems has predominantly focused on knowledge-intensive tasks (amayuelas-etal-2024-knowledge; yin_large_2023), with limited attention to reasoning tasks like mathematics. Therefore, we propose an unsolvable mathematical dataset for this gap. To ensure unsolvable problems under the same distribution of solvable data sources, we construct the unsolvable dataset $\mathcal{D}_{u}$ by rewriting solvable problems in $\mathcal{D}_{a}$ into unsolvable forms. For a problem $\boldsymbol{x}_{i}$ with $k$ necessary mathematical conditions, a unique solution $\boldsymbol{\hat{y}}_{i}$ is derived by reasoning using all conditions $\{{\boldsymbol{c}_{i}^{1}},\dots,{\boldsymbol{c}_{i}^{k}}\}\in\boldsymbol{x}_{i}$, as in Fig. 2 (a). Any alteration to one condition may result in the loss of the unique solution (fan2025missingpremiseexacerbatesoverthinking). Accordingly, we employ two schemes to synthesize unsolvable problems:
i) Removal: As in Fig. 2 (b), by removing one premise $\boldsymbol{c}_{i}^{2}$}, the original reasoning steps are prevented, rendering the problem unsolvable.
ii) Contradiction: As in Fig. 2 (c), introducing a condition $\boldsymbol{\tilde{c}}_{i}^{1}$} that contradicts $\boldsymbol{c}^{1}_{i}$} leads to logical inconsistency of $\boldsymbol{x}_{i}$, making it impossible to derive an exact solution of the problem.
Simply modifying a condition may result in false unsolvable cases where some rewritten problems have solutions that differs from the original, rather than genuinely unsolvable. Therefore, we propose a rigorous workflow in Sec. 3.3.